Let MC = 8+6Q and let MR = AR = P be fixed at P=20. What is the profit maximising output
for this firm?
1
Expert's answer
2011-11-18T10:54:00-0500
Let MC = 8+6Q and let MR = AR = P be fixed at P=20. What is the profit maximizing output for this firm?
The equilibrium condition is: MR=MC. In our case we have perfectly competitive firm: P=MC=AR.
MC=MR=P
8+6Q=20 Q=2;